OpenStudy (anonymous):

Is each situation a linear or a nonlinear model? Column A Column B 1. The amount of money a contractor charges for materials and labor is shown by the equation c = 550 + 30h, where c is the amount the contractor charges and h is the number of hours she works. 2. A car’s value is given by the equation y = 20,000(0.12)x, where y is the value of the car and x is the number of years since the car was purchased. 3. The number of times a person’s heart beats in several minutes is given by the equation b = 75m, where b is the number of heartbeats and m is the number of minutes.

OpenStudy (anonymous):

@gerryliyana @abdullah1995 @Jordan8901

OpenStudy (anonymous):

c = 550 + 30h and y = 20,000(0.12)x and b = 75m are linear model So 1, 2, and 3 are linear model

OpenStudy (anonymous):

so they are all linear?

OpenStudy (anonymous):

Of course

OpenStudy (anonymous):

ok thank you thats what i thought it was just was not for sure thank you

OpenStudy (anonymous):

you are welcome @NunChuckBoy3101 for sure i am going to draw it to the graph

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

2 is non :(

OpenStudy (anonymous):

For model 1

OpenStudy (anonymous):

for model 2

OpenStudy (anonymous):

For model 3

OpenStudy (anonymous):

2 i wrong it was non linear but thank you any way

OpenStudy (anonymous):

all graph show that they are linear

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

it said non so :(

OpenStudy (anonymous):

ok i'm sorry for that

OpenStudy (anonymous):

its ok